OpenStudy (anonymous):

The expression shown represents a function: x^2+4x-12 For what positive value of x will the graph of the function cross the x-axis? For what negative value of x will the graph of the function cross the x-axis?

OpenStudy (muzzack):

Step by step solution : Step 1 : Simplify x2+4x - 12 Trying to factor by splitting the middle term 1.1 Factoring x^2+4x-12 The first term is, x^2 its coefficient is 1 . The middle term is, +4x its coefficient is 4 . The last term, "the constant", is -12 Step-1 : Multiply the coefficient of the first term by the constant 1 • -12 = -12 Step-2 : Find two factors of -12 whose sum equals the coefficient of the middle term, which is 4 . -12 + 1 = -11 -6 + 2 = -4 -4 + 3 = -1 -3 + 4 = 1 -2 + 6 = 4 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, -2 and 6 x^2 - 2x + 6x - 12 Step-4 : Add up the first 2 terms, pulling out like factors : x • (x-2) Add up the last 2 terms, pulling out common factors : 6 • (x-2) Now add up the four terms of step 3 : (x+6) • (x-2) Which is the desired factorization Final result : (x + 6) • (x - 2)

OpenStudy (anonymous):

My sheet says the answer must be a single, regular number like: 2, 6, -5, etc. Would I multiply 6 and -2?

OpenStudy (muzzack):

yes

OpenStudy (anonymous):

So, is 6 the positive answer and -12 the negative answer?

OpenStudy (anonymous):

I mean -2

OpenStudy (muzzack):

yeah